OpenStudy (anonymous):

which of the following equations would graph a line parallel to 3y=2x-x+5? a.3y=x+1 b.y=x+1 c.y=2/3x+1 d.3/2y=x+1

OpenStudy (anonymous):

@humphriescody @walter20

OpenStudy (anonymous):

A 3y=x+1

OpenStudy (anonymous):

are u sure? @taylor1216

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

 yes, see they have the same slope and different intercepts meaning they are parallel

OpenStudy (anonymous):

can u help me with some more questions? @taylor1216

OpenStudy (anonymous):

Yeah. what are they?

OpenStudy (anonymous):

plans for a stadium are drawn on a coordinate grid. One wall lies on the line y=2x+3. A perpendicular wall passes through the points (4,-7). Write the equation of the line that contains the new wall.

OpenStudy (anonymous):

a.y=-1/2x-5 b.y=2x-15 c.y=1/2x-9 d.y=-2x=1

OpenStudy (anonymous):

@taylor1216 @humphriescody

OpenStudy (anonymous):

is it the same one

OpenStudy (anonymous):

what do u mean? @humphriescody

OpenStudy (anonymous):

@midhun.madhu1987 mighi be abel to help

OpenStudy (anonymous):

y=-1/2x-5

OpenStudy (midhun.madhu1987):

Parallel lines have equal slopes

OpenStudy (midhun.madhu1987):